Suppose you are testing the claim that the proportion of Democrats who are happy with their nominating process is the same as the proportion of Republicans who are happy with theirs. If you calculate a test statistic of z​ = 1.96, what is the​ P-value of your​ test? Use a level of significance​ = 0.05 A. .0250 B. .0500 C. .0020 D. .0040
Answer #1

Solution :

This is the two tailed test .

Test statistic = z =1.96

P(z > 1.96) = 1 - P(z <1.96 ) = 1-0.9750=0.025

P-value = 2*0.025=0.05

correct option b

P-value =

= 0.05

P-value <

Reject the null hypothesis .